OpenStudy (anonymous):

I just need someone to let me know if I am right?! help!? Let g(x) be the reflection of f(x) = 3x + 2 in the x-axis. What is a function rule for g(x)? my answer is b am i correct? g(x) = 3x – 2 g(x) = –3x – 2 g(x) = 3x + 2 g(x) = –3x + 2

jimthompson5910 (jim_thompson5910):

good you got it f(x) = 3x+2 g(x) = -f(x) g(x) = -(3x+2) g(x) = -3x-2

OpenStudy (anonymous):

ah thank you for letting me know!

jimthompson5910 (jim_thompson5910):

np